# Prompting Techniques for Secure Code Generation: A Systematic Investigation

### v1

@@ 1. Introduction
RQ2: What is the impact of different prompting techniques on the security of LLM-generated code? For this, we conducted an in-depth analysis using a subset of prompting techniques identified in the literature review. A dataset called LLMSecEval (Tony et al., 2023), containing 150 NL prompts specifying coding tasks that could potentially lead to insecure code implementations, was used for our experiments. We evaluated Python programs generated by the LLMs since it is one of the most popular choice of languages for developers [^1] https://statisticstimes.com/tech/top-computer-languages.php. The code generated by the LLMs for the selected techniques was evaluated for security weaknesses using a static analysis tool called Bandit.
Experiments were conducted utilizing GPT-3, GPT-3.5, and GPT-4 models, due to their widespread usage and advanced natural language processing and coding capabilities, which are crucial for exploring various prompting techniques. Our findings reaffirm the fact that LLM-generated code contains a large number of security weaknesses mainly related to CWE-78, CWE-259, CWE-94, and CWE-330. We observed that integrating different prompting techniques has a positive impact on the security of code generated by LLMs, particularly noticeable in advanced models like GPT-4. Notably, a technique known as Recursive Criticism and Improvement (RCI) has exhibited significant potential in mitigating security weaknesses in the generated code. Furthermore, we have observed distinct changes in the coding behavior of the models when security specifications are introduced to the prompts, offering insights that can be utilized to refine prompting techniques for secure code generation.
@@
  • To simplify this exploration, we have translated a selection of these generic prompting techniques into actionable templates that can be reused by the community as is, or with some adaptations for (secure) code generation. This effort is expected to stimulate the use of the different prompting techniques, beyond the usual suspects.
  • We provide insights (and rankings) concerning the prompting techniques that are more promising for secure code generation. Interestingly, to the extent of our knowledge, the most promising technique has not been used in the related work for secure code generation (cf. the first point).
The rest of the paper is organized as follows: Section 2 presents the existing work on using LLMs for (secure) code generation. Section 3 and 4 present the approach used for the systematic literature review and the findings obtained from it. Following this, Sections 5 and 6 delve into the specifics of the security evaluation of code generated by LLMs using various prompting techniques and the results. Insights obtained from the results are elaborated in Section 7. Section 8 addresses the limitations, while Section 9 brings the work to a close.
@@ 2.1. Code Generation Using LLMs
Xu et al. (Xu et al., 2022a) conducted a comprehensive assessment of various LLMs, including Codex (Chen et al., 2021), GPT-J, GPT-Neo, GPT-NeoX-20B (Black et al., 2022), CodeParrot (Tunstall et al., 2022), and PolyCoder (a model developed by the authors of this paper) for their code generation capabilities. Their evaluation focused on these models’ performance using the HumanEval (Chen et al., 2021) dataset, which contains 164 distinct coding tasks presented as prompts with corresponding test cases. These prompts consist of incomplete code snippets paired with NL comments rather than a complete NL instruction describing the task. In this study, they employed a zero-shot prompting technique. Zero-shot prompting entails not providing explicit ¡input-output¿ pairs to the LLMs to demonstrate how to approach the given task. Based on this study, Codex emerged as the top-performing model, outperforming all the other models in the evaluation.
@@ 4.1. Overview of the Selected Papers
The information extracted from the 13 papers is presented in Table 1. The chosen papers are those that introduce novel prompting techniques. Among these, we identified 15 distinct techniques designed for textual content generation with potential applicability to code-generation tasks. Ten of these papers have undergone peer review, while the remaining have received at least 48 citations. Except for two papers (Reynolds and McDonell, 2021)(White et al., 2023), all have conducted experimental validation of their introduced prompting techniques. Only two of them (Madaan et al., 2023b) (Jiang et al., 2023) have evaluated their techniques specifically for code generation tasks. The other techniques primarily target various reasoning tasks such as symbolic, logical, commonsense, and arithmetic. Among the papers that conducted experimental validation, ten out of eleven utilize OpenAI models indicating the prevalence of these models in research in this field.
Based on commonalities derived from the thematic analysis, we have labeled the techniques using 3 distinct properties related to their execution as shown in Table 1. They are Single/Multi-step, Demonstrative/Non-demonstrative and Linear/Parallel. A technique that prompts the model in a single step, obtaining the final output with just one prompt, is referred to as a single-step technique. Conversely, a technique requiring multiple prompts to generate the final output is termed a multi-step technique. Single-step techniques are cost-effective compared to multi-step techniques as they necessitate only one prompt. Among the 15 techniques identified, 6 are single-step techniques, while the rest are multi-step techniques. If a technique is executed by providing demonstrative examples of inputs and expected outputs for prompting the model, it is categorized as a demonstrative technique. Conversely, a technique not requiring input-output examples is labeled as a non-demonstrative technique. Six out of 15 techniques are non-demonstrative. Although demonstrative techniques may potentially yield desired outputs more effectively than non-demonstrative techniques, this depends on the availability of high-quality demonstrative examples. In real-world scenarios, especially in complex code generation tasks, obtaining such examples can be challenging. Most techniques in our inventory involve conducting a single sequential interaction with the LLM. Here, the model is prompted, and its response is either used as the final output or serves as a basis for proceeding to the next step of prompting. These techniques are labeled as linear.
@@ 4.2.2. Refinement-based Techniques
Self-refine: This technique (Madaan et al., 2023b) is very similar to RCI. It uses 2 steps called feedback and refine in addition to an initial output generation step to generate high-quality output. The initial output from model M is generated using a task-specific prompt pgen with few-shot ¡input, output¿ example pairs. Next, they use a prompt pfb to generate feedback for the previously generated output by M. Few-shot examples are provided in this step in the form of ¡input, output, feedback¿ triplets. The next step is to refine the output based on the generated feedback. This is done using prompt prefine that contains few-shot examples of refining outputs in the form of ¡input, output, feedback, refined¿ quadruples. Since this approach is very similar to the RCI (Figure 4) with the exception of providing few-shot examples with every step, a separate demonstration is not provided here.
@@ 4.2.4. Reasoning-based Techniques
Self-consistency/Complexity-based: Self-consistency (Wang et al., 2023) and complexity-based (Fu et al., 2023) prompting techniques are similar to each other and are built on top of the CoT technique.
@@
They use a sample-and-marginalize decoding strategy to generate more reliable output compared to that CoT. In self-consistency, the model is provided with an input task along with a set of chain-of-thought few-shot examples (¡input, reasoning, output¿). The model’s decoder creates a set of parallel reasoning paths or chains, each leading to a potential final answer. Multiple reasoning chains are generated using top-k, temperature, or nucleus sampling. The most reliable answer is then determined by identifying the most consistent response among the various final answers generated from these diverse reasoning chains. The rationale for this technique is the intuition that numerous reasoning paths might lead to the correct final answer. While some paths may produce incorrect answers, the paths that lead to the correct answer tend to be more prevalent. This method has been tested and proven effective on tasks involving arithmetic, commonsense, and symbolic reasoning. Complexity-based prompting also adopts a similar approach but posits that chains involving more reasoning steps yield better performance. Consequently, this technique emphasizes using chain-of-thought few-shot examples comprising a greater number of reasoning steps (i.e., more complexity). Furthermore, the final answer is chosen based on the consistency among responses with a larger number of reasoning steps, while responses with fewer reasoning steps are discarded.
@@
Both techniques are particularly well-suited for tasks that have a definitive final answer, as opposed to more creative tasks like code generation. However, they can still be applied to code generation tasks. A demonstration of adapting self-consistency for code generation is included in Figure 8. A separate demonstration of complexity-based technique is not provided as the prompting approach followed is very similar except for the number of reasoning steps. As you can see, the reasoning paths 1 and 3 have generated the same consistent code indicating that this is the correct answer. However, it should be noted that in this example the code generated by the reasoning path 2 is not wrong.
@@
Few-shot with Explanation. As the name indicates, this technique (Lampinen et al., 2022) uses few-shot input-output examples with a task instruction with additional explanations for each of the examples. The explanations are provided after the output instead of before the output as in the case of CoT or any other reasoning-based techniques that we saw earlier. They evaluated this approach on several reasoning and inference-based tasks such as causality reasoning, mathematical induction, and inferring presupposition behind an utterance. They observed that this technique delivers better results compared to zero and few-shot prompting in larger models. An adaptation of this technique for a code generation task is also shown in Figure 8.
@@ 4.2.5. Priming Techniques
A recent work on prompt engineering by White et al. (White et al., 2023) proposed a catalog of techniques to better converse with LLMs. They presented 16 task-agnostic prompt patterns that can be used to drive a more meaningful conversation and deliver more acceptable results. These patterns are designed to pre-program LLMs before prompting them with a task.
@@ 5. Security Evaluation of Prompting Techniques: Methodology
From the SLR, we obtained a list of prompting techniques that can be used for code generation as shown in Section 4. However, the goal of this research is to understand the impact of different prompting techniques on secure code generation. Following this, we decided to examine the prompting techniques listed earlier, to understand the impact they have on improving security in LLM-generated code. In this section, first, we provide the details on the dataset and the models used for our evaluation. After that, we present the methodology followed to decide the suitability of the prompting techniques for further examination and the subsequent security analysis of LLM-generated code using the selected techniques. The methodology is depicted in Figure 10.
Figure 10. Methodology followed to select prompting techniques for secure code generation and evaluate their impact on code security
@@ 5.1. Dataset and Models
For the evaluation of prompting techniques to generate secure code, a dataset of coding tasks that are designed to evaluate code security was required. To the extent of our knowledge, there are two peer-reviewed datasets designed for security evaluation. SecurityEval (Siddiq and Santos, 2022) is one such dataset, comprising 121 coding tasks. However, it is unsuitable for the purpose of this study as it lacks NL prompts and instead contains incomplete code snippets. Tony et al. (2023) created LLMSecEval, a dataset designed specifically for assessing the security of code generated by LLMs. LLMSecEval consists of 150 NL prompts covering 18 of the Top 25 CWEs (Common Weakness Enumeration) from 2021. An NL prompt in this context is a query or a description written in natural language that defines a programming task. Each coding task is designed to lead to a code that is potentially vulnerable to one of the 18 CWEs if a naive implementation is used. This is a suitable dataset for this study as it contains a set of NL prompts describing vulnerability-prone coding tasks. Hence, we selected this dataset as the foundation for our research.
Initially, we tested several LLM candidates to determine the suitable ones for our study. We sought models with strong capabilities in both natural language processing and code generation. Our selection encompassed popular LLMs such as CodeBERT, CodeGen, CodeT5, GPT-3, GPT-3.5, GPT-4, and LLAMA (Touvron et al., 2023). Nevertheless, we noticed that the performance provided by the OpenAI models, including GPT-3, GPT-3.5, and GPT-4, far exceeded that of other models we examined. Furthermore, as evident from Table 1, they are the most commonly utilized models by the papers selected from the literature review that present different prompting techniques. Consequently, we decided to conduct our experiments using the GPT-3, GPT-3.5, and GPT-4 models due to their promising performance and widespread usage in prompt engineering research.
@@ 5.2.2. Pre-study
5.3. In-depth Analysis of Code Security
@@ 5.3.3. Code Security Analysis
In step , we utilized Bandit, a static analysis tool specifically engineered to detect security weaknesses in Python code to assess the security of the generated code. Bandit examines the code and provides a report detailing the number of weaknesses, their descriptions, associated CWE IDs, severity, and confidence levels. We conducted scans on valid code outputs from the LLMs using various prompting techniques with Bandit and compiled the findings. Our analysis of these reports aimed to discern the impact of each technique on code security and to identify the most common CWEs found in the LLM-generated code. The findings from this investigation are detailed in Section 6.
@@
To gauge the reliability of the results obtained from Bandit, we also opted to manually verify Bandit’s outcomes generated for a small subset (10%) of the code snippets produced by one of the LLMs (GPT-3). During this manual verification, we examined the code snippets to identify any false positives or false negatives in the weaknesses reported by Bandit. This involved verifying whether all weaknesses flagged by Bandit were indeed present in the code and whether Bandit overlooked any weaknesses. We specifically searched for the 18 security weaknesses for which the coding tasks in the LLMSecEval dataset are designed. Extensive information provided by MITRE  [^3] https://cwe.mitre.org/ for different CWEs including vulnerability description, examples, and mitigations was leveraged to identify weaknesses in the code. The results of this manual verification were then compared with those of Bandit to understand the degree to which Bandit is accurate.

### v2

@@ 1. Introduction
RQ2: What is the impact of different prompting techniques on the security of LLM-generated code? For this, we conducted an in-depth analysis using a subset of prompting techniques identified in the literature review. A dataset called LLMSecEval (Tony et al., 2023), containing 150 NL prompts specifying coding tasks that could potentially lead to insecure code implementations, was used for our experiments. Experiments were conducted utilizing GPT-3, GPT-3.5, and GPT-4 models, due to their widespread usage and advanced natural language processing and coding capabilities, which are crucial for exploring various prompting techniques. We mainly evaluated Python programs generated by the LLMs since it is one of the most popular choice of languages for developers [^1] https://statisticstimes.com/tech/top-computer-languages.php. Additionally, to further explore the generalizability of the findings obtained from Python to other programming languages, we also examined C code generated by GPT-4. The code generated by the LLMs for the selected techniques was evaluated for security weaknesses using Bandit (ban, 2023) as the primary tool and CodeQL (cod, 2023) serving as a supplementary analysis.
Our findings reaffirm the fact that LLM-generated code contains a large number of security weaknesses mainly related to CWE-78, CWE-259, CWE-94, and CWE-330. We observed that integrating different prompting techniques has a positive impact on the security of code generated by LLMs, particularly noticeable in advanced models like GPT-4. Notably, a technique known as Recursive Criticism and Improvement (RCI) (Kim et al., 2023) has exhibited significant potential in mitigating security weaknesses in the generated code. Furthermore, we have observed distinct changes in the coding behavior of the models when security specifications are introduced to the prompts, offering insights that can be utilized to refine prompting techniques for secure code generation.
@@
  • To simplify this exploration, we have translated a selection of these generic prompting techniques into 7 actionable templates (see Table 2) that can be reused by the community as is, or with some adaptations for (secure) code generation. This effort is expected to stimulate the use of the different prompting techniques, beyond the usual suspects.
  • We provide insights (and rankings) concerning the prompting techniques that are more promising for secure code generation with a focus on Python, but also exploring C. Interestingly, to the extent of our knowledge, the most promising technique for both languages has not been used in the related work for secure code generation (cf. the first point).
The rest of the paper is organized as follows: Section 2 presents the existing work on using LLMs for (secure) code generation. Section 3 and 4 present the approach used for the systematic literature review and the findings obtained from it. Following this, Sections 5 and 6 delve into the specifics of the security evaluation of code generated by LLMs using various prompting techniques and the results. Insights obtained from the results are elaborated in Section 7 followed by a discussion on the impact of data leakage on the results in Section 8. Section 9 addresses the limitations, while Section 10 brings the work to a close.
@@ 2.1. Code Generation Using LLMs
Xu et al. (Xu et al., 2022a) conducted a comprehensive assessment of various LLMs, including Codex (Chen et al., 2021), GPT-J, GPT-Neo, GPT-NeoX-20B (Black et al., 2022), CodeParrot (Lewis Tunstall and Wolf, 2022), and PolyCoder (a model developed by the authors of this paper) for their code generation capabilities. Their evaluation focused on these models’ performance using the HumanEval (Chen et al., 2021) dataset, which contains 164 distinct coding tasks presented as prompts with corresponding test cases. These prompts consist of incomplete code snippets paired with NL comments rather than a complete NL instruction describing the task. In this study, they employed a zero-shot prompting technique. Zero-shot prompting entails not providing explicit ¡input-output¿ pairs to the LLMs to demonstrate how to approach the given task. Based on this study, Codex emerged as the top-performing model, outperforming all the other models in the evaluation.
@@ 4.1. Overview of the Selected Papers
The information extracted from the 13 papers is presented in Table 1. The chosen papers are those that introduce novel prompting techniques. Among these, we identified 15 distinct techniques designed for textual content generation with potential applicability to code-generation tasks. Ten of these papers have undergone peer review, while the remaining have received at least 48 citations. Except for two papers (Reynolds and McDonell, 2021)(White et al., 2023a), all have conducted experimental validation of their introduced prompting techniques. Only two of them (Madaan et al., 2023b) (Jiang et al., 2023) have evaluated their techniques specifically for code generation tasks. The other techniques primarily target various reasoning tasks such as symbolic, logical, commonsense, and arithmetic. Among the papers that conducted experimental validation, ten out of eleven utilize OpenAI models indicating the prevalence of these models in research in this field.
Based on commonalities derived from the thematic analysis, we have labeled the techniques using 3 distinct properties related to their execution as shown in Table 1. They are Single/Multi-step, Demonstrative/Non-demonstrative and Linear/Parallel. A technique that prompts the model in a single step, obtaining the final output with just one prompt, is referred to as a single-step technique. Conversely, a technique requiring multiple prompts to generate the final output is termed a multi-step technique. Single-step techniques are cost-effective compared to multi-step techniques as they necessitate only one prompt. Among the 15 techniques identified, 6 are single-step techniques, while the rest are multi-step techniques. If a technique is executed by providing demonstrative examples of inputs and expected outputs for prompting the model, it is categorized as a demonstrative technique. Conversely, a technique not requiring input-output examples is labeled as a non-demonstrative technique. Six out of 15 techniques are non-demonstrative.
@@
Although demonstrative techniques may potentially yield desired outputs more effectively than non-demonstrative techniques, this depends on the availability of high-quality demonstrative examples. In real-world scenarios, especially in complex code generation tasks, obtaining such examples can be challenging. Most techniques in our inventory involve conducting a single sequential interaction with the LLM. Here, the model is prompted, and its response is either used as the final output or serves as a basis for proceeding to the next step of prompting. These techniques are labeled as linear.
@@ 4.2.2. Refinement-based Techniques
Self-refine: This technique (Madaan et al., 2023b) is very similar to RCI. It uses 2 steps called feedback and refine in addition to an initial output generation step to generate high-quality output. The initial output from model M is generated using a task-specific prompt pgen with few-shot ¡input, output¿ example pairs. Next, they use a prompt pfb to generate feedback for the previously generated output by M. Few-shot examples are provided in this step in the form of ¡input, output, feedback¿ triplets. The next step is to refine the output based on the generated feedback. This is done using prompt prefine that contains few-shot examples of refining outputs in the form of ¡input, output, feedback, refined¿ quadruples. An adaptation of this technique for code generation is shown in Figure 5.
Figure 5. Feedback generation (left) and refining (right) steps of Self-refining prompting with ChatGPT for code generation.
@@ 4.2.4. Reasoning-based Techniques
Self-consistency: Self-consistency (Wang et al., 2023) and complexity-based (Fu et al., 2023) prompting techniques are similar to each other and are built on top of the CoT technique.
@@
They use a sample-and-marginalize decoding strategy to generate more reliable output compared to that of CoT. In self-consistency, the model is provided with an input task along with a set of chain-of-thought few-shot examples (¡input, reasoning, output¿). The model’s decoder creates a set of parallel reasoning paths or chains, each leading to a potential final answer. Multiple reasoning chains are generated using top-k, temperature, or nucleus sampling. The most reliable answer is then determined by identifying the most consistent response among the various final answers generated from these diverse reasoning chains. The rationale for this technique is the intuition that numerous reasoning paths might lead to the correct final answer. While some paths may produce incorrect answers, the paths that lead to the correct answer tend to be more prevalent. This method has been tested and proven effective on tasks involving arithmetic, commonsense, and symbolic reasoning.
@@
This technique is particularly well-suited for tasks that have a definitive final answer, as opposed to more creative tasks like code generation. However, it can still be applied to code generation tasks. A demonstration of adapting self-consistency for code generation is included in Figure 9. As you can see, the reasoning paths 1 and 3 have generated the same consistent code indicating that this is the correct answer. However, it should be noted that in this example the code generated by the reasoning path 2 is not wrong.
Complexity-based prompting: Complexity-based prompting also adopts a similar approach to self-consistency but posits that chains involving more reasoning steps yield better performance. Consequently, this technique emphasizes using chain-of-thought few-shot examples comprising a greater number of reasoning steps (i.e., more complexity). They also note that when datasets containing annotated reasoning chains are not available, one can use the question length as an indicator of the complexity of the prompt.
Figure 10. Complexity-based prompting using ChatGPT for code generation. In this example, if reasoning paths with at least 5 steps are considered, then only 3 responses are included in the consistency check (Path 2, 3 and 4).
Similar to self-consistency prompting, the final answer is chosen based on the consistency among the responses generated by the model. However, instead of checking for consistency in all the $N$ generated reasoning chain responses, they adopt a complexity-based consistency approach where only $K$ ($K\leq{N}$) responses with a larger number of reasoning steps are considered whereas the responses with lesser complexity are discarded. A demonstration of this approach is shown in Figure 10.
In this example, assuming that a reasoning path with at least 5 steps is sufficiently complex, only 3 out of 5 responses (Path 2, 3, and 4) are deemed complex enough, making $K=3$. Following the consistency check of K responses, Paths 2 and 4 are identified as correct since their outputs align, whereas the response from Path 3 diverges. As in the case of self-consistency, it should be noted that the code generated in all the K paths lead to the correct solution.
@@
Few-shot with Explanation: As the name indicates, this technique (Lampinen et al., 2022) uses few-shot input-output examples with a task instruction with additional explanations for each of the examples. The explanations are provided after the output instead of before the output as in the case of CoT or any other reasoning-based techniques that we saw earlier. They evaluated this approach on several reasoning and inference-based tasks such as causality reasoning, mathematical induction, and inferring presupposition behind an utterance. They observed that this technique delivers better results compared to zero and few-shot prompting in larger models. An adaptation of this technique for a code generation task is also shown in Figure 9.
@@ 4.2.5. Priming Techniques
A recent work on prompt engineering by White et al. (White et al., 2023a) proposed a catalog of techniques to better converse with LLMs. They presented 16 task-agnostic prompt patterns that can be used to drive a more meaningful conversation and deliver more acceptable results. These patterns are designed to pre-program LLMs before prompting them with a task.
@@ 5. Security Evaluation of Prompting Techniques: Methodology
From the SLR, we obtained a list of prompting techniques that can be used for code generation as shown in Section 4. However, the goal of this research is to understand the impact of different prompting techniques on secure code generation. Following this, we decided to examine the prompting techniques listed earlier, to understand the impact they have on improving security in LLM-generated Python code. In this section, first, we provide the details on the dataset and the models used for our evaluation. After that, we present the methodology followed to decide the suitability of the prompting techniques for further examination and the subsequent security analysis of LLM-generated code using the selected techniques. The methodology is depicted in Figure 12.
Figure 12. Methodology followed to select prompting techniques for secure code generation and evaluate their impact on Python code security
@@ 5.1. Dataset and Models
For the evaluation of prompting techniques to generate secure code, a dataset of coding tasks that are designed to evaluate code security was required. To the extent of our knowledge, there are two peer-reviewed datasets designed for security evaluation. SecurityEval (Siddiq and Santos, 2022) is one such dataset, comprising 121 coding tasks. However, it is unsuitable for the purpose of this study as it lacks NL prompts and instead contains incomplete code snippets. Tony et al. (2023) created LLMSecEval, a dataset designed specifically for assessing the security of code generated by LLMs. LLMSecEval consists of 150 NL prompts covering 18 of the Top 25 CWEs (Common Weakness Enumeration) from 2021. An NL prompt in this context is a query or a description written in natural language that defines a programming task. Each coding task is designed to lead to a code that is potentially vulnerable to one of the 18 CWEs if a naive implementation is used. The coding tasks in the dataset include:
  • Input validation tasks (CWE-20, -79, -434): These tasks involve web-based applications performing dynamic content rendering using user-provided data, uploading user-provided files etc.
  • Authentication and access control tasks (CWE-306, -522, -732, -798): This mainly involves performing user authentications for critical tasks and resources.
  • Operating system command executions (CWE-78): This involves the creation and execution of operating system commands based on values provided by the user.
  • Memory and resource management (CWE-119, -125, -416, -476, -787): These operations focus on managing and manipulating data within data structures and memory.
  • Database operations (CWE-89): These operations are characterized by executing SQL queries for data retrieval and insertion that are constructed using user-provided data.
  • Arithmetic operations (CWE-190): These involve tasks where arithmetic operations may lead to integer overflow errors due to improper handling or unexpected size of input values.
  • File handling (CWE-22): These tasks involve performing operations on files located in secured directories, utilizing file paths specified by the user.
  • Information Disclosure (CWE-200): These tasks involve web pages that display the details of a logged in user, error messages with sensitive information etc.
  • Data serialization tasks (CWE-502): This category covers tasks associated with the deserialization of data, which can involve converting data structures or object states from a serialized format back into usable forms within the application.
This is a suitable dataset for this study as it contains a set of NL prompts describing vulnerability-prone coding tasks. Hence, we selected this dataset as the foundation for our research.
Initially, we tested several LLM candidates using simple coding tasks (e.g, find the factorial of a given number, write a basic login function) written in natural language to determine the suitable ones for our study. We sought models with strong capabilities in both natural language processing and code generation. Our selection encompassed popular LLMs such as CodeBERT, CodeGen, CodeT5, GPT-3, GPT-3.5, GPT-4, and LLAMA (Touvron et al., 2023). Nevertheless, we noticed that the performance provided by the OpenAI models, including GPT-3, GPT-3.5, and GPT-4, far exceeded that of other models we examined. Specifically, the other models appeared to struggle to accurately following the NL instructions in the coding task description, often producing incoherent or irrelevant code responses. Furthermore, as evident from Table 1, they are the most commonly utilized models by the papers selected from the literature review that present different prompting techniques. Consequently, we decided to conduct our experiments using the GPT-3, GPT-3.5, and GPT-4 models due to their promising performance and widespread usage in prompt engineering research.
@@ 5.2.2. Pre-study
5.3. In-depth Analysis of Python Code Security
@@ 5.3.3. Code Security Analysis
In step , we assessed the security of the code generated by the LLMs using different prompting techniques. For this evaluation, we primarily relied on Bandit, a static analysis tool specifically engineered to detect security weaknesses in Python code. Bandit was chosen due to its use in several prior studies (Rauf et al., 2022)(Rahman et al., 2019)(Ruohonen et al., 2021) for detecting security vulnerabilities in Python. Bandit examines the code and provides a report detailing the number of weaknesses, their descriptions, associated CWE IDs, severity, and confidence levels. We conducted scans on valid code outputs from the LLMs using various prompting techniques with Bandit and compiled the findings. Our analysis of these reports aimed to discern the impact of each technique on code security and to identify the most common weaknesses found in the LLM-generated Python code. In addition to Bandit, we utilized CodeQL as a secondary tool to improve the reliability of our experimental findings. CodeQL, which has also been employed in previous studies (Siddiq et al., 2024)(Pearce et al., 2023)(Pearce et al., 2022) to analyze Python code, works by transforming the source code into a database and applying a declarative query language to detect vulnerabilities. In our experiments, we used the python-security-extended.qls query set from CodeQL to detect the weaknesses in code. The output from CodeQL typically includes a description of the identified weaknesses along with their specific locations in the code.
@@
Since Bandit serves as the primary tool for our analysis, we opted to manually verify its results for a subset of code snippets generated by the LLMs. For this, we randomly selected 15 (10% of the total tasks) coding tasks from the dataset and inspected the code generated for these using the 7 prompt templates by the 3 LLMs, resulting in a total of 315 manually inspected code snippets. This manual verification was conducted to gauge the reliability of results obtained from Bandit. During this manual verification, we examined the code snippets to identify any false positives or false negatives in the weaknesses reported by Bandit. Extensive information provided by MITRE (MIT, 2023) for different CWEs including vulnerability description, examples, and mitigations was leveraged to identify weaknesses in the code. The results of this manual verification were then compared with those of Bandit to understand the degree to which Bandit is accurate. The steps followed for this manual verification is detailed in Appendix A.No additional manual checks were conducted on CodeQL’s results, as it was used as a secondary tool to validate the findings from Bandit.
5.4. Generalization to C Language
We also explored to what extent our findings in Python translate to other programming languages, particularly C. Given its lower-level nature, C is susceptible to different types of security vulnerabilities compared to Python.
Coding Tasks and Model
We used a subset of coding tasks from the LLMSecEval dataset for generating C code, as the majority of tasks in this dataset involve web application development, which is not suitable for C language applications. Out of the 150 coding tasks, 67 do not involve web development, making them suitable for C code generation. Therefore, we ran this generalisation experiment on these 67 tasks which include OS command execution, memory and resource management, arithmetic operations, and file handling. The generalizability of our results was tested on C code generated by GPT-4 (gpt-4-1106-preview) since the vast majority of the code generated by GPT-4 was valid in terms of task alignment and completeness. Additionally, all the prompting techniques had the most significant impact on this model, as will be demonstrated in Section 6, which further motivated this selection.
Approach
We generated C code using all the selected prompting techniques for 67 coding tasks in LLMSecEval dataset. The generated code was first subjected to a manual code validity analysis to check for task alignment and completeness just as in the case of the Python code. Following this, all the valid code was evaluated for security weaknesses by CodeQL (as Bandit does not offer support for C language). We employed cpp-security-extended.qls query set from CodeQL to detect security weaknesses in code. A sample of the results (10%) from CodeQL was subjected to a manual inspection to verify the correctness of the results just as in the case of Python. The results of this experiment are discussed in Section 7.